Jump to main content.


Drake Chemical

EPA Weekly Situation Reports for September 1998

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Wednesday, September 30, 1998

FOR DATES: From 7:00 a.m. Wednesday, Sept. 23, 1998 to 7:00 a.m. Wednesday, Sept. 30, 1998.

The Thermal Destruction Facility (TDF) was on-line for 162.3 hours and processed 5638.5 tons of contaminated soil during the time frame of this report. The TDF was on-line 96.6 % of the time with an average hourly feed rate of 34.7 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated 145,479.0 tons of contaminated soil, which amounts to approximately 52.7 % of the overall contaminated soil on-site.

During the regular TDF operations, the off-line time was primarily attributed to the stoppage of feed during local lightning storm events. Feed was stopped as a precaution to possible power outages that are normally associated with thunderstorms. Additional off-line time was attributed to the stoppage of feed to perform regular calibrations to the TDF equipment and when the feed screws became bound with soil material.

Construction activities during the time frame of this report included TDF operations, wastewater management, feed preparation, debris separation, lime treatment of soils, ash handling, ash backfilling, dust control to prevent fugitive emissions, the placement odor suppressant foam on the excavation face, rock and concrete debris cleaning and backfilling.

The contractor continued to perform real time perimeter air monitoring and off-site air monitoring activities. MRI, OHM's subcontractor performing perimeter and off-site air sampling, temporarily shut down the real time air monitoring system on Wednesday, September 23, in order to perform preventative maintenance and leak checks on the heat trace lines. Real time air sampling was restarted on Thursday afternoon, once the testing and maintenance were completed. During the down time, perimeter site air was monitored with hand held devices by the contractor.

Excavation continued along the southern end of the excavation face and along the corridor sheetpiling. In addition, excavation of sediment basin 1 continued as part of the site grading plan within the exclusion zone. No drums were unearthed during the time frame of this report. Two roll-offs of PPE and non-hazardous debris were shipped off-site to a hazardous waste landfill.

The weekly progress meeting was held on Tuesday, September 29, 1998 with representatives from the USEPA, USACE, PADEP, and OHM in attendance.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, and rock backfilling. The TDF is tentatively scheduled to be shut down (hot hold) on October 6th and 7th in order to perform several maintenance tasks, including the replacement of the dual feed screws.

PREPARED BY: Steve Gibson for David Modricker, USACE


Thursday, September 24, 1998

FOR DATES: From 7:00 a.m. Wednesday, Sept. 16, 1998 to 7:00 a.m. Wednesday, Sept. 23, 1998.

The Thermal Destruction Facility (TDF) was on-line for 157.38 hours and processed 5669.7 tons of contaminated soil during the time frame of this report. The TDF was on-line 93.7 % of the time with an average hourly feed rate of 36.0 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated 140,205.8 tons of contaminated soil, which amounts to approximately 50.8 % of the overall contaminated soil on-site. The operations phase of the project was indicated to be 48.2 % complete at the end of this reporting period. OHM has indicated that the 50 % completion milestone should be attained within the next two weeks.

During the regular TDF operations, the off-line time was attributed to the stoppage of feed during daily calibrations of the TDF equipment and short term feed stoppages due to the dual feed screws becoming bound with wet soil feed materials. The contractor has indicated that the reduced feed rate (less than target) is a result of the soil materials binding to the dual feed screws. The soil that binds to the feed screws causes the screw motor to work harder, which could potentially trip the motor breaker and stop feed to the incinerator.

For this reporting period, the off-line time was primarily attributed to a planned TDF warm shutdown in order to perform maintenance to the feed screws and to perform the monthly calibrations to the TDF equipment. In addition to the monthly calibrations, the contractor performed the monthly TDF interlock checks to test the incinerator's automatic shut down systems. As part of the monthly interlock checks, the contractor initiated a planned Thermal Relief Valve (TRV) opening. The TRV was opened at 1030 hours on Thursday, September 17. Prior to the TRV opening, feed had been shut down for approximately 3 hours and the proper authorities were notified.

Construction activities during the time frame of this report included TDF operations, wastewater management, feed preparation, debris separation, wood debris shredding, lime treatment of soils, ash handling, ash backfilling, dust control to prevent fugitive emissions, the placement of Concover on recently graded areas of the stockpile and bottom ash stockpile, rock and concrete debris cleaning and backfilling. The contractor continued to perform real time perimeter air monitoring and off-site air monitoring activities.

Excavation was continued along the southern end of the excavation face and along the corridor sheetpiling. In addition to the routine mass excavation activities, sediment basin 1 was cleaned out and prepared for phase out with respect to the site grading plan within the exclusion zone. No drums were unearthed during the time frame of this report.

Related site activities during this reporting period include:

* The weekly progress meeting was held on Tuesday, September 22, 1998 with representatives from the USEPA, USACE, PADEP, and OHM in attendance.

* The quarterly stack testing for metals/particulate was completed on Wednesday, September 16, 1998.

* A representative for the AIR group toured the site to observe the completion of the stack testing activities.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, and rock backfilling.

PREPARED BY: Steve Gibson for David Modricker, USACE


Wednesday, September 16, 1998

FOR DATES: From 7:00 a.m. Wednesday, Sept. 9, 1998 to 7:00 a.m. Wednesday, Sept. 15, 1998.

The Thermal Destruction Facility (TDF) was on-line for 160.0 hours and processed 6419 tons of contaminated soil during the time frame of this report. The TDF was on-line 95.2 % of the time with an average hourly feed rate of 40.1 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated 133,124.7 tons of contaminated soil, which amounts to approximately 48.2 % of the overall contaminated soil on-site.

Off-line time is attributed primarily to the stoppage of feed during daily calibrations of the TDF equipment and to perform preventative maintenance to the drive chain that turns the rotary kiln. Feed was stopped for approximately 4.5 hours on the evening of Friday, September 11, 1998 in order to safely remove an injured worker from the feed preparation building.

As stated above, feed to the incinerator was stopped because an OHM employee was injured on Friday evening in the feed preparation building. The worker apparently tripped on the stairs of the weighbelt platform and fell onto the side of a debris bin, injuring his back. In order to avoid aggravation to a possible back injury, a local ambulance crew with HAZWOPER training arrived on-site to safely move the worker from the feed preparation building to the local emergency room. The worker was examined at Lock Haven Hospital and was released to return to work the same night.

Construction activities during the time frame of this report included TDF operations, wastewater management, feed preparation, debris separation, wood debris shredding, lime treatment of soils, ash handling, ash backfilling, dust control to prevent fugitive emissions, the placement of Concover on recently graded areas of the stockpile and bottom ash stockpile, rock and concrete debris cleaning and backfilling. The contractor continued to perform real time perimeter air monitoring and off-site air monitoring activities.

Excavation was continued at the southwest corner of the excavation. No drums were unearthed during the time frame of this report. Two roll-offs containing non-hazardous waste (i.e. scrap metal, PPE, and lime bag debris) were shipment off-site to a hazardous waste landfill.

Related site activities during this reporting period include:

* The weekly progress meeting was held on Tuesday, September 15, 1998 with representatives from the USEPA, USACE, PADEP, and OHM in attendance.

* Representatives from the USACE and OHM attended a Local Emergency Planning Committee (LEPC) meeting on Tuesday, September 15, 1998.

* Carmine Leone and Benny Hom of the USACE North Atlantic Division (NAD) arrived on-site to tour the Drake Chemical Superfund Site and several other local ACOE projects.

* Two runs of the quarterly stack testing for metals and particulates were completed during this reporting period. The third and final run of this quarter's stack testing was completed Wednesday afternoon.

* A representative for the AIR group toured the site to observe stack testing and site activities.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, and rock backfilling.

PREPARED BY: Steve Gibson for David Modricker, USACE


Wednesday, September 9, 1998

FOR DATES: From 7:00 a.m. Wednesday, Sept 2, 1998 to 7:00 a.m. Wednesday, Sept. 9, 1998.

The Thermal Destruction Facility (TDF) was on-line for 163.5 hours and processed 6592.3 tons of contaminated soil during the time frame of this report. The TDF was on-line 97.3 % of the time with an average hourly feed rate of 40.3 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated 126,617 tons of contaminated soil, which amounts to approximately 46 % of the overall contaminated soil on-site.

Off-line time is attributed primarily to the stoppage of feed during local lightning storm events and during daily calibrations of the TDF equipment.

Construction activities during the time frame of this report included TDF operations, wastewater management, feed preparation, debris separation, wood debris shredding, lime treatment of soils, ash handling, ash backfilling, dust control to prevent fugitive emissions, the placement of Concover on recently graded areas of the stockpile, and rock and concrete debris handling, cleaning and backfilling. The contractor continued to perform real time perimeter air monitoring and off-site air monitoring activities.

No excavation was performed on the south side of the contaminated soil stockpile during the time frame of this report. Scrap metal and lime bag debris was washed and placed in rolloffs for shipment off-site to a hazardous waste landfill.

Related site activities during this reporting period include the weekly progress meeting that was held on Tuesday, September 8, 1998 with representatives from the USEPA, USACE, PADEP, and OHM in attendance.

Representatives from the USACE North Atlantic Division (NAD) will be touring the site early next week.

Stack testing for metals and particulates is scheduled for September 14 through 16, 1998.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, and rock backfilling.

PREPARED BY: Steve Gibson for David Modricker, USACE


Wednesday, September 2, 1998

FOR DATES: From 7:00 a.m. Wednesday, August 26, 1998 to 7:00 a.m. Wednesday, September 2, 1998.

The Thermal Destruction Facility (TDF) was on-line for 166.9 hours and processed 6715 tons of contaminated soil during the time frame of this report. The TDF was on-line 99.3 % of the time with an average hourly feed rate of 40.2 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated 118,172 tons of contaminated soil, which amounts to approximately 43 % of the overall contaminated soil on-site.

Nineteen new hires for OHM arrived at the Drake Site on Monday, August 31, 1998, after successfully completing the OSHA 40-hour HAZWOPER training. The new hires consist of remedial technicians, equipment operators, and TDF mechanics. The new employees are currently involved in additional 'on the job' training at the site to become familiar with remedial work and safety precautions specific to the Drake Chemical Superfund Site.

Construction activities during the time frame of this report included wastewater management, excavation within the exclusion zone, feed preparation, debris separation, wood debris chipping, lime treatment of soils, ash handling, ash backfilling, dust control to prevent fugitive emissions, rock and concrete debris handling, cleaning and backfilling. The contractor performed the monthly survey of the bottom ash stockpile as an ash volume indicator.

The contractor continued to perform real time perimeter air monitoring and off-site air monitoring activities. The time weighted average monthly perimeter air sampling event began at the end of this reporting period.

No drums were excavated during the time frame of this report. Demolition debris consisting of wood, metal, and concrete continued to be unearthed and separated in the exclusion zone.

Related site activities include:

* The weekly progress meeting was held on Tuesday, September 1, 1998 with representatives from the USEPA, USACE, PADEP, and OHM in attendance.

* Representatives of the Drake Cleanup Team met with the Lock Haven Environmental Advisory Committee, and later with the Clinton County Farm Bureau on Tuesday, September 1, 1998.

* Stack testing for metals and particulates is scheduled for September 14 through 16, 1998.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning & backfilling, drum handling, and drum bulking.

PREPARED BY: Steve Gibson for David Modricker, USACE

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.